OpenStudy (yinshy):

Describe the addition pattern and find the next two terms. 25, 28, 31, 34, 37, . . . A. Three is added to a term to get the next term in the pattern. The next two terms are 41 and 44 B. Three is added to a term to get the next term in the pattern. The next two terms are 40 and 43 C. Four is added to a term to get the next term in the pattern. The next two terms are 41 and 45 D. Four is added to a term to get the next term in the pattern. The next two terms are 40 and 44

OpenStudy (sacredtyras):

It's A.

OpenStudy (caozeyuan):

28-25=31-28=3, so...

OpenStudy (caozeyuan):

@sacredtyras , should b b